# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v4

@@ 4.2 Latent-level sampling
The complexity of modeling distributions directly at the output level can be mitigated by using generative models, which often rely on simpler, lower-dimensional latent variables $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to instead learn the approximate through
@@
In contrast, the mutual information term here encapsulates the aleatoric uncertainty, while the expected entropy reflects the epistemic uncertainty. The key overarching contribution of these methods lies in the conditioning of generative models on the input image. This section briefly introduces Generative Adversarial Networks (Section 4.2.1), Variational Autoencoders (Section 4.2.2), and Denoising Diffusion Probabilistic Models (Section 4.2.3), and outlines how these architectures are commonly adapted for probabilistic segmentation.
@@ 4.2.1 Generative Adversarial Networks (GANs)
$\mathcal{L}_{\text{adv}}=-\mathbb{E}_{p_{\mathcal{D}}}\mathbb{E}_{p_{\mathbf{Z}}}[\log D_{\bm{\psi}}(G_{\bm{\phi}}(F_{\bm{\theta}}(\mathbf{x}),\mathbf{z}),\mathbf{x}),\mathbf{x})],$ (16)
@@ 4.2.2 Variational Autoencoders (VAEs)
$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p(\mathbf{Y}|\mathbf{X})\,]$
$=\mathbb{E}_{q{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{p(\mathbf{z}|\mathbf{Y},\mathbf{X})}\right]$
$\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\operatorname{KL}[\,q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\,].$ (20)
The first term in Equation (20) represents the reconstruction cost of the decoder, subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$. The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and prior density. As a consequence of the mean-field approximation, all involved densities are modeled by axis-aligned Gaussian densities and amortized through neural networks, parameterized by $\bm{\phi}$, $\bm{\theta}$ and $\bm{\psi}$. The predictive distribution after observing dataset $\mathcal{D}$ is then obtained as
@@
$\log p(\mathbf{x})=\log p_{\*Z}(\mathbf{z}_{0})-\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|,$ (22)
[…]
@@
$\hskip 10.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}|\mathbf{X})\,]-\mathbb{E}_{q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})}\left[\,\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|\,\right],$ (23)
where the objective consists of a reconstruction term, sample-based KL-divergence and a likelihood correction term for the change in probability density induced by the NF.
Bhat et al. (2023; 2022a) compare this approach with other parameterizations of the latent space, including a mixture of Gaussians and low-rank approximation of the full covariance matrix. Valiuddin et al. (2024b) show that the latent space can converge to contain non-informative latent dimensions, undermining the capabilities of the latent-variable approach, generally referred to as mode or posterior collapse (Chen et al., 2016; Zhao et al., 2017). Their proposition considers the alternative formulation of the ELBO, specified as
@@
$\hskip 35.0pt-\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}).$
The proposed novel objective maximizes the contribution of the (expected) mutual information between latent and output variables, i.e. the aleatoric uncertainty term of the predictive entropy in Equation (13). This can be rewritten into the objective
@@
Hyperparameter $\lambda$ is the Lagrange multiplier update through the Exponential Moving Average of the reconstruction, which is constrained to reach target value $\kappa$, empirically set beforehand to an appropriate value. Finally, online negative hard mining is used to only backpropagate 2% of the worst performing pixels, which are stochastically picked with the Gumbel-SoftMax trick (Jang et al., 2016; Huijben et al., 2022). A similar architecture known as PHiSeg (Baumgartner et al., 2019) takes a similar approach to the HPU-Net. However, PHiSeg places residual connections between latent vectors across decompositions rather than in the convolutional layers. Furthermore, deep supervision at each decomposition scale to enforce the disentanglement between latent variables.
@@ 4.2.3 Denoising Diffusion Probabilistic Models (DDPMs)
5 Epistemic uncertainty
The key distinction between epistemic and aleatoric uncertainty lies in their origins: epistemic uncertainty arises from model ignorance, whereas aleatoric uncertainty reflects inherent statistical variability in the data. Consequently, unlike aleatoric uncertainty, epistemic uncertainty should not only be quantified but ideally also minimized. Epistemic uncertainty can be further categorized into two distinct types (Hüllermeier & Waegeman, 2021). The first type pertains uncertainty related to the capacity of the model. For example, under-parameterized models or approximate model posteriors, that can become too stringent to appropriately resemble the true posterior. The ambiguity on the best parameters given the limited capacity induces uncertainty in the learning process, which is also known as model uncertainty. Nevertheless, given the complexity of contemporary parameter-intensive CNNs, the model uncertainty is often assumed to be negligible. A more significant contribution to the epistemic uncertainty is due to the limited data availability, known as approximation uncertainty, and can often be reduced by collecting more data.
@@ 5.1 Variational Inference
Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$, to approximate posterior $p(\bm{\theta}|\mathbf{y},\mathbf{x})$. Then, we can employ Variational Inference (VI) w.r.t. to the parameters, by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior as
$\bm{\eta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\operatorname{KL}\left[\,q(\bm{\theta}|\bm{\eta})\,||\,p(\bm{\theta}|\mathbf{y},\mathbf{x})\,\right]$
@@
where the parameter-independent terms are constant and therefore excluded from notation. A popular choice for the approximated variational posterior is the Gaussian distribution, i.e. a mean $\mu$ and covariance $\sigma$ parameter for each element of the convolutional kernel, usually with zero-mean unitary Gaussian prior densities. However, the priors can be also learned through Empirical Bayes (Bishop, 1995). Furthermore, backpropagation is possible with the reparameterization trick (Kingma & Welling, 2013) and within this context, the procedure is referred to as Bayes by Backprop (BBB) (Blundell et al., 2015) and has later been improved with the Local Reparameterization trick (Kingma et al., 2015). In addition to approximating the posterior, sampling from it yields multiple parameter permutations, effectively enriching the model’s hypothesis space, a phenomenon also referred to as model combination (Minka, 2000; Clarke, 2003; Lakshminarayanan, 2016). In most cases, VI is not performed explicitly. Instead, simpler techniques are employed, with the resulting parameter permutations hypothesized to serve as a proxy. We discuss two such methods: Monte Carlo Dropout (Section 5.1.1) and Ensembling (Section 5.1.2). See Figure 11 for an illustration of each method. See Figure 11 for an illustration of each method. All of these methods have been widely applied to segmentation. However, unlike aleatoric uncertainty modeling, their implementations are often more generic and not specifically tailored to segmentation tasks. Therefore, we provide only a brief overview here, followed by a more in-depth discussion in later sections related to various applications.
@@ 5.1.1 Monte Carlo Dropout
with probability $p$ and vector $\mathbf{n}$ operating element-wise on the parameters. It has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process (Gal & Ghahramani, 2016). In this manner, such a method is able to provide multimodal estimates of the model uncertainty. As noted by Gal et al. (2017a), the variance in the model output is primarily governed by the magnitudes of the weights rather than the dropout rate $p$, which is usually optimized through grid search or simply set to $0.5$. Hence, the authors propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, to increase the influence of $p$. As the name suggests, a continuous approximation to the discrete distribution is used, known as the Concrete distribution (Maddison et al., 2016; Jang et al., 2016), to enable path-wise derivatives through $p$. Also, a generalization of MC dropout has been proposed, where the weights of the network, rather that its hidden output, are set to zero (Mobiny et al., 2021).
@@ 5.2 Laplace approximation
It can be shown that the intractable posterior can be estimated as
@@ 5.3 Test-time augmentation
An image $\mathbf{X}$ can be understood as a one-of-many visual representations of the object of interest. For example, systematic noise, translation or rotation result in many realistic variations that approximately retain image semantics. Hence, data augmentation (Krizhevsky et al., 2012) has been used at test time (explaining the name test-time augmentation, or TTA), argued to obtain uncertainty estimates by efficiently exploring the locality of the likelihood function (Ayhan & Berens, 2022). By randomly augmenting input images with invertible transformation $\tilde{\mathbf{x}}=T_{\zeta}(\mathbf{x})$, with transformation parameters $\zeta$, a prediction is obtained with $\tilde{\mathbf{y}}=f_{\bm{\theta}}(\tilde{\mathbf{x}})$ and can then be inverted through $\mathbf{y}=T_{\zeta}^{-1}(\tilde{\mathbf{y}})$. Repeatedly performing this procedure results in a set of segmentation masks, which serve as an estimate of $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$. In segmentation literature, it has been hypothesized that TTA encapsualtes the aleatoric uncertainty in a better way (Wang et al., 2019a; Ayhan & Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Whitbread & Jenkinson, 2022; Roshanzamir et al., 2023). However, literature opposing this also exist and even suggests that TTA enables the modeling of epistemic uncertainty (Hu et al., 2022). Kahl et al. (2024) draw parallels with BNNs and demonstrate this through decomposing the predictive uncertainty as
@@ 6.1 Observer variability
DDPM Chen et al. (2023); Zbinden et al. (2023); Rahman et al. (2023)
@@ 6.2 Active Learning
Selecting the most informative samples for training contains two elements. Firstly, the uncertainty needs to be quantified. This is usually the predictive entropy Kasarla et al. (2019); Shen et al. (2021); Xie et al. (2022) or variance Ozdemir et al. (2021), but in some instances the epistemic uncertainty is explicitly estimated through Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011; Ma et al., 2024b; Shen et al., 2021). Simply stated, the reduction in posterior entropy indicates the informative of a data sample. In turn, this can also be formulated by the reduction in predictive entropy in the dual formulation. We can formally denote this as
@@
$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-H[\mathbf{y}|\bm{\theta},\mathbf{x}^{*},\mathcal{D}]$
@@
Note the similarity between the mutual information obtained from the predictive entropy decomposition in Equation (3). Secondly, to evaluate whether the active learning strategy is beneficial, the performance is mostly measured subject to increasingly stringent budget requirements (i.e. a percentage of the original dataset) Yang et al. (2017); Sourati et al. (2018); Kasarla et al. (2019); Mahapatra et al. (2018); Sinha et al. (2019); Siddiqui et al. (2020); Li & Alstrøm (2020); Kim et al. (2021); Shen et al. (2021); Burmeister et al. (2022); Ma et al. (2024b). Kahl et al. (2024) argue that Active Learning should be evaluated on image-level as humans dont annotate on pxiel-level. Assuming a saturated performance on in-distribution data (training cycle $t_{1}$), the author propose to measure the improved relative performance, $C$, subject to a second training sample ($t_{2}$) from selected samples
@@
Since the nature of this problem involves identifying and reducing model ignorance, the quantification of epistemic uncertainty is most appropriate. Hence, as can be seen in Table 2, most of the literature use MC dropout of ensembles. Furhtermore, using the plain SoftMax activation has also been commonly used. Often, solely relying on the most uncertain samples can result in samples with limited diversity and therefore samples selected on the representativeness Ozdemir et al. (2021); Shen et al. (2021); Wu et al. (2021); Mahapatra et al. (2018); Sinha et al. (2019). Another crucial modeling element pertains the aggregation of the pixel-level uncertainties. Therefore, closeness to boundary Gorriz et al. (2017); Kasarla et al. (2019); Ma et al. (2024b) or specific regions Gaillochet et al. (2023); Wu et al. (2021); Kasarla et al. (2019) are additionally incorporated when combining the uncertainty values across pixels. Siddiqui et al. (2020) determine the entropy through estimating the entropy across viewpoints in multi-view datasets. Gaillochet et al. (2023) argue the superiority of selecting samples based on batch-level uncertainty.

### v5

@@ 4.2 Latent-level sampling
The limitation of modeling intricate, complex distributions directly at the output level can be mitigated by using generative models, which often rely on simpler, lower-dimensional latent variables $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to instead learn the approximate through
@@
In contrast to Equation (3), the mutual information term here encapsulates the aleatoric uncertainty, while the expected entropy reflects the epistemic uncertainty. Also, note that the mutual information is between the output and latent variables, rather than the parameters of the model. The key overarching contribution of latent-level sampling methods lies in the conditioning of generative models on the input image. This section briefly introduces Generative Adversarial Networks (Section 4.2.1), Variational Autoencoders (Section 4.2.2), and Denoising Diffusion Probabilistic Models (Section 4.2.3), and outlines how these architectures are commonly adapted for probabilistic segmentation.
@@ 4.2.1 Generative Adversarial Networks (GANs)
$\mathcal{L}_{\text{adv}}=-\mathbb{E}_{p_{\mathcal{D}}}\mathbb{E}_{p_{\mathbf{Z}}}[\log D_{\bm{\psi}}(G_{\bm{\phi}}(F_{\bm{\theta}}(\mathbf{x}),\mathbf{z}),\mathbf{x})],$ (16)
@@ 4.2.2 Variational Autoencoders (VAEs)
$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\log\int q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\mathrm{d}\mathbf{z}$ (20a)
$\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\right]$ (20b)
$=\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\operatorname{KL}[\,q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\,],$ (20c)
where Jensen’s inequality justifies moving the expectation outside the logarithm. The first term in Equation (20) represents the reconstruction cost of the decoder, subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$. The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and prior density. As a consequence of the mean-field approximation, all involved densities are modeled by axis-aligned Gaussian densities and amortized through neural networks, parameterized by $\bm{\phi}$, $\bm{\theta}$ and $\bm{\psi}$. The predictive distribution after observing dataset $\mathcal{D}$ is then obtained as
@@
$\log p(\mathbf{x})=\log p_{\mathbf{Z}}(\mathbf{z}_{0})-\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|,$ (22)
where $\mathbf{z}_{k}$ and $\mathbf{z}_{k-1}$ are intermediate variables from intermediate densities and $\mathbf{z}_{0}=\mathbf{f}^{-1}(\mathbf{x})$. Equation (22) can be substituted in the conditional ELBO objective in Equation (20) to obtain
@@
$\hskip 10.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}_{k}|\mathbf{X})\,]-\mathbb{E}_{q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})}\left[\,\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|\,\right],$ (23)
where the objective consists of a reconstruction term, sample-based KL-divergence and a likelihood correction term for the change in probability density induced by the NF. Bhat et al. (2023; 2022a) compare this approach with other parameterizations of the latent space, including a mixture of Gaussians and low-rank approximation of the full covariance matrix. Valiuddin et al. (2024b) show that the latent space can converge to contain non-informative latent dimensions, undermining the capabilities of the latent-variable approach, generally referred to as mode or posterior collapse (Chen et al., 2016; Zhao et al., 2017). Their proposition considers the alternative formulation of the ELBO, specified as
@@
$\hskip 35.0pt-\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}).$
The proposed novel objective maximizes the contribution of the (expected) mutual information between latent and output variables, i.e. the aleatoric uncertainty term of the predictive entropy in Equation (13). This enables the introduction of the updated objective
@@
Hyperparameter $\lambda$ is the Lagrange multiplier update through the Exponential Moving Average of the reconstruction, which is constrained to reach target value $\kappa$, empirically set beforehand to an appropriate value. Finally, online negative hard mining is used to only backpropagate 2% of the worst performing pixels, which are stochastically picked with the Gumbel-SoftMax trick (Jang et al., 2016; Huijben et al., 2022). Furthermore, PHiSeg (Baumgartner et al., 2019) takes a similar approach to the HPU-Net. However, PHiSeg places residual connections between latent vectors across decompositions rather than in the convolutional layers. Furthermore, deep supervision at each decomposition scale to enforce the disentanglement between latent variables.
@@ 4.2.3 Denoising Diffusion Probabilistic Models (DDPMs)
5 Parameter modeling
The key distinction between epistemic and aleatoric uncertainty lies in their origins: epistemic uncertainty arises from model ignorance (i.e. related to the parameters), whereas aleatoric uncertainty reflects inherent statistical variability in the data. Consequently, unlike aleatoric uncertainty, epistemic uncertainty should not only be quantified but ideally also minimized. Epistemic uncertainty can be further categorized into two distinct types (Hüllermeier & Waegeman, 2021). The first type pertains uncertainty related to the capacity of the model. For example, under-parameterized models or approximate model posteriors, that can become too stringent to appropriately resemble the true posterior. The ambiguity on the best parameters given the limited capacity induces uncertainty in the learning process, which is also known as model uncertainty. Nevertheless, given the complexity of contemporary parameter-intensive CNNs, the model uncertainty is often assumed to be negligible. A more significant contribution to the epistemic uncertainty is due to the limited data availability, known as approximation uncertainty, and can often be reduced by collecting more data.
@@ 5.1 Variational Inference
Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$ (e.g. a gaussian with $\bm{\eta}=\{\mu,\sigma\}$), to approximate posterior $p(\bm{\theta}|\mathbf{y},\mathbf{x})$. Then, we can employ Variational Inference (VI) w.r.t. to the parameters, by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior as
$\bm{\eta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\operatorname{KL}\left[\,q(\bm{\theta}|\bm{\eta})\,||\,p(\bm{\theta}|\mathcal{D})\,\right]$
@@
where the parameter-independent terms are constant and therefore excluded from notation. In the case of a deterministic encoder, i.e. $q(\bm{\theta}|\bm{\eta})=\delta(\bm{\theta}-\bm{\theta}^{*})$, the formulation collapses to the MAP estimate in Equation (5). A popular choice for the approximated variational posterior is the Gaussian distribution, i.e. a mean $\mu$ and covariance $\sigma$ parameter for each element of the convolutional kernel, usually with zero-mean unitary Gaussian prior densities. However, the priors can be also learned through Empirical Bayes (Bishop, 1995). Furthermore, backpropagation is possible with the reparameterization trick (Kingma & Welling, 2013) and within this context, the procedure is referred to as Bayes by Backprop (BBB) (Blundell et al., 2015) and has later been improved with the Local Reparameterization trick (Kingma et al., 2015). In addition to approximating the posterior, sampling from it yields multiple parameter permutations, effectively enriching the model’s hypothesis space, a phenomenon also referred to as model combination (Minka, 2000; Clarke, 2003; Lakshminarayanan, 2016). In most cases, VI is not performed explicitly. Instead, simpler techniques are employed, with the resulting parameter permutations hypothesized to serve as a proxy. We discuss two such methods: Monte Carlo Dropout (Section 5.1.1) and Ensembling (Section 5.1.2). See Figure 11 for an illustration of each method. All of these methods have been widely applied to segmentation. However, unlike aleatoric uncertainty modeling, their implementations are often more generic and not specifically tailored to segmentation tasks. Therefore, we provide only a brief overview here, followed by a more in-depth discussion in later sections related to various applications.
@@ 5.1.1 Monte Carlo Dropout
with probability $p$ and vector $\mathbf{n}$ operating element-wise on the parameters. It has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process (Gal & Ghahramani, 2016). In this manner, such a method is able to provide multimodal estimates of the model uncertainty. As noted by Gal et al. (2017a), the variance in the model output is primarily governed by the magnitudes of the weights rather than the dropout rate $p$. The dropout rate is usually optimized through grid search or simply set to $p=0.5$. The relationship between $p$ and the magnitude of the model weights has also be exploited to probabilistically prune neural networks Gonzalez-Carabarin et al. (2022). Gal et al. (2017a) propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, to increase the influence of $p$. As the name suggests, a continuous approximation to the discrete distribution is used, known as the Concrete distribution (Maddison et al., 2016; Jang et al., 2016), to enable path-wise derivatives through $p$. Also, a generalization of MC dropout has been proposed, where the weights of the network, rather that its hidden output, are set to zero (Mobiny et al., 2021).
@@ 5.2 Laplace approximation
It can be shown that the intractable posterior in Equation (2) can be estimated as
@@ 5.3 Test-time augmentation
An image $\mathbf{X}$ can be understood as a one-of-many visual representations of the object of interest. For example, systematic noise, translation or rotation result in many realistic variations that approximately retain image semantics. Hence, data augmentation (Krizhevsky et al., 2012) has been used at test time (explaining the name test-time augmentation, or TTA), argued to obtain uncertainty estimates by efficiently exploring the locality of the likelihood function (Ayhan & Berens, 2022). By randomly augmenting input images with invertible transformation $\tilde{\mathbf{x}}=T_{\zeta}(\mathbf{x})$, with transformation parameters $\zeta$, a prediction is obtained with $\tilde{\mathbf{y}}=f_{\bm{\theta}}(\tilde{\mathbf{x}})$ and can then be inverted through $\mathbf{y}=T_{\zeta}^{-1}(\tilde{\mathbf{y}})$. Repeatedly performing this procedure results in a set of segmentation masks, which serve as an estimate of $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$. In segmentation literature, it has been hypothesized that TTA encapsualtes the aleatoric uncertainty in a better way (Wang et al., 2019a; Ayhan & Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Whitbread & Jenkinson, 2022; Roshanzamir et al., 2023). However, literature opposing this also exist and even suggests that TTA enables the modeling of epistemic uncertainty (Hu et al., 2022). Kahl et al. (2024) draw parallels with BNNs and demonstrate this through the (by now) familiar decomposition of the predictive uncertainty
@@ 6.1 Observer variability
DDPM Chen et al. (2023a); Zbinden et al. (2023); Rahman et al. (2023)
@@ 6.2 Active Learning
Selecting the most informative samples for training contains two elements. Firstly, the uncertainty needs to be quantified. This is usually the predictive entropy Kasarla et al. (2019); Shen et al. (2021); Xie et al. (2022) or variance Ozdemir et al. (2021), but in some instances the epistemic uncertainty is explicitly estimated through Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011; Ma et al., 2024b; Shen et al., 2021). Simply stated, the reduction in posterior entropy indicates the informativeness of a data sample. In turn, this can also be formulated by the reduction in predictive entropy in the dual formulation. We can formally denote this as
@@
$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-H[\mathbf{y}|\bm{\theta},\mathbf{x}^{*},\mathcal{D}],$
@@
Note the similarity between the mutual information $I$ obtained from the predictive entropy decomposition in Equation (3), where both are the objective of epistemic uncertainty modeling. Secondly, to evaluate whether the active learning strategy is beneficial, the performance is mostly measured subject to increasingly stringent budget requirements (i.e. a percentage of the original dataset) Yang et al. (2017); Sourati et al. (2018); Kasarla et al. (2019); Mahapatra et al. (2018); Sinha et al. (2019); Siddiqui et al. (2020); Li & Alstrøm (2020); Kim et al. (2021); Shen et al. (2021); Burmeister et al. (2022); Ma et al. (2024b). Kahl et al. (2024) argue that Active Learning should be evaluated on image-level as humans dont annotate on pxiel-level. Assuming a saturated performance on in-distribution data (training cycle $t_{1}$), the author propose to measure the improved relative performance, $C$, subject to a second training sample ($t_{2}$) from selected samples
@@
Since the nature of this problem involves identifying and reducing model ignorance, the quantification of epistemic uncertainty is most appropriate. Hence, as can be seen in Table 2, most of the literature use MC dropout of ensembles. Furthermore, using the plain SoftMax activation has also been commonly used. Often, solely relying on the most uncertain samples can result in samples with limited diversity and therefore samples selected on the representativeness Ozdemir et al. (2021); Shen et al. (2021); Wu et al. (2021); Mahapatra et al. (2018); Sinha et al. (2019). Another crucial modeling element pertains the aggregation of the pixel-level uncertainties. Therefore, closeness to boundary Gorriz et al. (2017); Kasarla et al. (2019); Ma et al. (2024b) or specific regions Gaillochet et al. (2023); Wu et al. (2021); Kasarla et al. (2019) are additionally incorporated when combining the uncertainty values across pixels. Siddiqui et al. (2020) determine the entropy through estimating the entropy across viewpoints in multi-view datasets. Gaillochet et al. (2023) argue the superiority of selecting samples based on batch-level uncertainty.